OpenStudy (anonymous):

Convert the polar equation to rectangular from: r= 3sec theta / 5 sec theta +5

OpenStudy (michele_laino):

please note that I can write your equation as below: \[r=\frac{ 3 }{ 5 }\frac{ \cos(\theta+5) }{ \cos \theta } \]

OpenStudy (anonymous):

Okay, that makes sense.

OpenStudy (michele_laino):

so we have: \[\frac{ 5 }{ 3 }r=\cos 5-\tan \theta \sin 5\] do you agree? @IronWalrus

OpenStudy (anonymous):

I don't understand how you got the right side of the equal sign. Can you explain that?

OpenStudy (michele_laino):

because: \[\frac{ \cos(\theta+5) }{ \cos \theta }=\frac{ \cos \theta \cos 5-\sin \theta \sin 5 }{ \cos \theta }\] \[=\cos5-\tan \theta \sin 5\]

OpenStudy (michele_laino):

do you agree? @IronWalrus